In a pipe of 400mm and 900m length an oil of specific gravity 0.75 is flowing at the rate of 0.5Rm³/s find

1) Heat lost due to friction and

2) power required viscosity of oil as 0.3 Stokes.

Take kinematic viscosity of oil as 0.3stroke.


1
Expert's answer
2021-06-10T10:23:48-0400

1)

"V=\pi d^2 l,"

"v=\frac Vt,"

"t=\frac Vv=\frac{\pi d^2 l}{v}=\frac{3.14\cdot 0.4^2\cdot 900}{0.5}=15~min,"

"\Delta Q=\rho \eta gt=0.75\cdot 0.3\cdot 10\cdot 900=2.02~kJ,"

2)

"P=\frac{\Delta Q}{t}=\frac{2020}{900}=2.25~W."